Find the eccentricity, foci and centre of the conic 4x^2 + 9y^2= 36. Draw a rough sketch of it.
1
Expert's answer
2020-10-26T16:42:39-0400

"\displaystyle
4x^2 + 9y^2 = 36\\

\frac{x^2}{9} + \frac{y^2}{4} = 1\\


\frac{(x - 0)^2}{9} + \frac{(y - 0)^2}{4} = 1\\


\textsf{The equation describes an ellipse}\\
\textsf{with the origin as its centre.}\\

a^2 = 9 \\

b^2 = 4\\


c^2 = a^2 - b^2 = 9 - 4 = 5\\

c = \pm\sqrt{5}\\


\therefore \textsf{The focus is}\, (\sqrt{5}, 0), (-\sqrt{5}, 0).\\


\textsf{Eccentricity}\, = \frac{c}{a} = \frac{\sqrt{5}}{3}\\"